OpenStudy (anonymous):

What is the major axis of the equation shown below? (x - 2)^2 / 49 + (y + 5)^2 / 36 = 1 Will give fans and medals.

OpenStudy (misty1212):

Hi!!

OpenStudy (misty1212):

this is the standard form of a certain shape do you know what it is?

OpenStudy (anonymous):

Sorry, I had forgotten to specify.

OpenStudy (misty1212):

i am asking if you know what it is parabola hyperbola ellipse

OpenStudy (anonymous):

I think it's an ellipse. However, I would truly appreciate it if you would help me find the major axis.

OpenStudy (misty1212):

yes it is an ellipse

OpenStudy (misty1212):

standard form is \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\]

OpenStudy (misty1212):

where the center is \((h,k)\) what is the center here?

OpenStudy (anonymous):

(2, -5)

OpenStudy (misty1212):

ok and next, which one is it?

OpenStudy (anonymous):

The one on the left.

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

we know the center is \((2,-5)\) and \(a^2=49\) so \(a=7\) and \(b^2=36\) so \(b=6\) what is left ?

OpenStudy (misty1212):

ooh the length of the major axis?

OpenStudy (anonymous):

Yes, please.

OpenStudy (misty1212):

ok you are going to be annoyed at how easy this one is we got \(a=7\) right?

OpenStudy (misty1212):

it is \(2a=14\)

OpenStudy (anonymous):

So, it's just 14?

OpenStudy (misty1212):

we can find the endpoints if you like but yes, it is just 14

OpenStudy (anonymous):

Man, that WAS easy. Thank you.

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (misty1212):

btw the length of the minor axis is .... 12

OpenStudy (anonymous):

Thank you for that too.